khozynhodges68:

Find the length of the hypotenuse of an isosceles right triangle whose legs are 1 unit in length.

justus:

Hello. Since it's an isosceles triangle it means any two sides are equal H^2=1^2+1^2 H=√2

justus:

So the right answer is √2 units.